# Online Multi-Task Offloading for Semantic-Aware Edge Computing Systems

### v2

@@ II-B Transmission Model
$t_{n}=\frac{\rho_{n}s_{n}\mu_{n}^{p}}{R_{n}},$ (2)
where $s_{n}$ is the normalized transmission data from UEn to ES, and $p>0$ is a constant parameter. $\rho_{n}$ is an offload indicator, where $\rho_{n}=1$ represents offloading tasks to ES, and $\rho_{n}=0$ denotes local computation. Compared to the amount of data required for task offloading, the computation results typically involve a much smaller data size [39, 40], therefore, we only consider the upload rate and latency.
@@
TABLE I: Notations and Symbols
Notation Explanation
@@
II-C Computation Model
@@
$t_{n}^{U}=\begin{cases}\frac{(1-\rho_{n})l_{n}^{U}}{c_{n}},\ \text{if}\ \mu_{n}=1,\\ \frac{(1-\rho_{n})l_{n}^{U}}{c_{n}}+\frac{\rho_{n}l_{n}^{SE}}{\mu_{n}^{q}c_{n}},\ \mu_{n}^{min}\leq\mu_{n}<1,\end{cases}$ (3)
where $q>0$ is a constant parameter, $l_{n}^{U}$ is the Graphic Processing Unit (GPU) consumption of the current task computed by UEn, and $l_{n}^{SE}$ is the basic GPU consumption of semantic extraction. $l_{n}^{U}$ and $l_{n}^{SE}$ are related to the corresponding task $\mathcal{M}$. $k>0$ is a constant parameter. $c_{n}=\alpha_{ma}N_{U}f_{n}$ is the GPU computation capability of UEn. $\alpha_{ma}=2$ corresponds to the floating-point operations (FLOPs) per cycle per core (FLOPs/cycle/core), where each instruction executes a fused multiply-add (FMA) operation. $N_{u}$ represents the number of CUDA cores available on UEn. $f_{n}$ denotes the GPU clock frequency of UEn. When tasks are offloaded to ES, the computation time of tasks on the ES can be expressed as
@@
II-D Energy Consumption Model
@@
$E^{S}=\kappa^{S}t^{S}f_{s}^{3},$ (7)
where $\kappa^{S}$ is the energy coefficient of ES. Therefore, the energy consumption of UEn’s current task $q_{n}^{\mathcal{M}}$ can be expressed as
@@
II-E Multi-Task Offloading Design and Problem Formulation
@@
$\varepsilon_{n}^{\mathcal{M}}=f^{\mathcal{M}}\left(s_{n},\nu_{n}\right),$ (9)
where $\varepsilon_{n}^{\mathcal{M}}$ is the accuracy of task $\mathcal{M}$, and $f^{\mathcal{M}}(\cdot)$ is the pretrained semantic transmission model for task $\mathcal{M}$, and $\nu_{n}$ is SNR. Once the model training is completed, we can directly calculate the number of FLOPs required by the UE’s semantic encoder to determine the basic GPU consumption $l_{n}^{SE}$ for semantic extraction. Similarly, we compute the number of FLOPs required by the ES’s semantic decoder to determine the GPU consumption $l_{n}^{S}$ for the ES. To evaluate the performance of offloading tasks for different UEs, it is essential to define a unified QoE standard.
@@
The performance of task execution is associated with computation time, energy consumption, and task accuracy. However, defining a fair and reasonable QoE is challenging. When tasks of different types are unevenly distributed among different UEs, directly calculating the weighted sum of energy consumption, computation time, and task accuracy is unfair. Typically, text tasks have lower computation time and energy consumption than image tasks, while image tasks have lower computation time and energy consumption than VQA tasks. Additionally, the accuracies of different tasks are usually incomparable. Inspired by [41, 42, 43], we propose a QoE calculation method that can fairly compare task execution performance across different UEs:
@@
where $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ are the weights of computation time, energy consumption, and task accuracy, respectively. The sum of $\omega_{t}$, $\omega_{e}$ and $\omega_{a}$ are equal to 1. $Q_{n}$ is the task length of UEn. $t^{l}_{n}=l_{n}^{U}/c_{n}$ denotes the task execution delay by UEn. As shown in Fig. 3, we use the logistic/sigmoid function and its symmetrical counterpart for QoE calculations. Corresponding latency score $G^{t}_{\mathcal{M}}$ and energy consumption score $G^{e}_{\mathcal{M}}$ are calculated using the symmetric logistic function, while accuracy score $G^{a}_{\mathcal{M}}$ is measured with the logistic function. The offset $x_{0}$ represents local execution latency, energy consumption, or accuracy. Specifically, $E_{n}^{l}=\kappa^{U}t_{n}^{l}(f_{n})^{3}$ denotes the task execution energy consumption by UEn. $\varepsilon_{n}^{l}$ denotes the task execution accuracy by UEn. That is, we normalize the QoE using the latency, energy consumption, and accuracy of tasks executed locally by the UE.
@@
[…]
@@
$\sum_{k=1}^{\mathcal{K}}x_{k,n}=1,\quad\forall n\in\mathcal{N},$ (11d)
$\sum_{n=1}^{\mathcal{N}}x_{k,n}=1,\quad\forall k\in\mathcal{K},$ (11e)
@@
where $\{\boldsymbol{\rho},\boldsymbol{p},\boldsymbol{f},\boldsymbol{x},\boldsymbol{\mu}\}=\{\rho_{n},p_{n},f_{n},x_{k,n},\mu_{n}\},\forall{n}\in\mathcal{N},\forall{k}\in\mathcal{K}$. $\boldsymbol{x}\in\mathbb{R}^{K\times N}$ denote the collection of all sub-channel allocation indicators $x_{k,n}$. $x_{k,n}=1$ indicates UEn is assigned to sub-channel $k$, $x_{k,n}=0$ otherwise. In (P0), (11b) and (11c) represent the offloading choice and channel selection constraints, respectively. Constraints (11d) and (11e) indicate that each UE is assigned to one sub-channel and each sub-channel is occupied by one UE. $p_{max}$ in (11g) is the maximum transmit power constraint, $f_{max}$ in (11h) is the maximum GPU frequency of UEn, $t_{max}$ in (11i) is the maximum execute latency constraint, $E_{n}^{max}$ in (11j) is the battery capacity of UEn, and $\varepsilon_{min}$ in (11k) is the minimum task accuracy requirement, which is dependent on the task type $\mathcal{M}$ of the current task.
III Multi-Agent Proximal Policy Optimization for Resource Allocation
The binary offloading variable in Constraint (11b) introduces non-convexity, and the QoE is characterized by a non-convex logistic function, rendering the problem P0 a non-convex MINLP problem. To solve the non-convex MINLP problem P0, we propose an advanced semantic-aware MAPPO algorithm, i.e., an actor-critic-based reinforcement learning algorithm, which is suitable for both discrete and continuous action spaces. To utilize the semantic-aware MAPPO algorithm, we reformulate problem P0 as an MDP and detail the training process of the MAPPO.
III-A Problem Reformulation Based on MDP
@@
A 4-tuple MDP $\langle\mathcal{S},\mathcal{A},\mathcal{R},\gamma\rangle$ is used to describe the interaction between the ES and the wireless network, where $\mathcal{S}$, $\mathcal{A}$, $\mathcal{R}$ and $\gamma\in[0,1)$ denote the state space, action space, reward function, and the discount factor, respectively. As agents, UEs need to collect observable states to input into the policy network and obtain their actions. After executing actions, UEs update their observable states and the policy network based on corresponding rewards to improve the output of actions. Specifically, the observable state of UEn can be represented as $o_{n}=\{g_{k,n},l_{n}^{U},l_{n}^{S}\}\in\mathcal{S}$. The action space of UEn refers to the variables that need to be optimized, which can be expressed as $a_{n}=\{\rho_{n},p_{n},f_{n},\mu_{n},x_{k,n}\}\in\mathcal{A}$. UEs need to continuously optimize offloading strategies, transmit power, local computation frequency, semantic compression ratio, and channel assignment to purchase better QoE. Drawing from the optimization problem P0, it is evident that the reward function needs to focus on the summation QoE of all UEs. In addition to QoE, the reward function includes penalty terms for latency and task accuracy, expressed as
[…]
@@
Therefore, the total reward at the $t$th training timestep is defined as $r_{t}=\sum^{N}_{n=1}r_{n}$. The cumulative long-term reward accounts for the impact of the current state and action on the expected future rewards:
@@
III-B Semantic-aware MAPPO Algorithm
@@
$J^{\boldsymbol{\theta}^{\prime}}(\boldsymbol{\theta})=\mathbb{E}_{\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\sim\pi_{\boldsymbol{\theta}^{\prime}}}\left[\frac{p_{\boldsymbol{\theta}}\left(\boldsymbol{a}_{t}\mid\boldsymbol{s}_{t}\right)}{p_{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{a}_{t}\mid\boldsymbol{s}_{t}\right)}A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\right],$ (17)
@@
$A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=Q_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)-V_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t}\right).$ (18)
The state-action value function $Q_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)$ is represented as
$Q_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=\mathbb{E}_{\boldsymbol{a}_{t}\sim\pi_{\boldsymbol{\theta^{\prime}}},\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}\right],$ (19)
@@
$V_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t}\right)=\mathbb{E}_{\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}\right].$ (20)
It is noted that $Q_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)$ can also be expressed by temporal difference (TD) form as $Q_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=r_{t}+\gamma V_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t+1}\right)$. Therefore, (18) can be reformulated as
$A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=r_{t}+\gamma V_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t+1}\right)-V_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t}\right).$ (21)
@@
$A^{GAE}(\boldsymbol{s}_{t},\boldsymbol{a}_{t})=\sum_{l=0}^{T-t}(\lambda\gamma)^{l}\delta_{t+l},$ (22)
where $\delta_{t}=A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)$. $\lambda\in(0,1]$ is a discount factor. Note that (22) aims to balance the variance and bias of the advantage estimates by leveraging information from multiple future steps while discounting their influence.
@@
If there are no constraints on the policies $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}^{\prime}}$, significant fluctuations may occur during policy updates. This is due to the limited sampling, which fails to accurately represent the probability distributions described in Lemma 1. To constrain the divergence between policies $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}^{\prime}}$, $J^{CLIP}(\boldsymbol{\theta})$ is employed. It is defined as
$J^{CLIP}(\boldsymbol{\theta})=L_{t}^{\mathrm{CLIP}}$ (23)
$=\mathbb{E}_{t}\left[\min\left(r_{t}(\boldsymbol{\theta})A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right),g\left(\epsilon,A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\right)\right)\right],$
@@
$g(\epsilon,$ $A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right))=$ (24)
$\begin{cases}(1+\epsilon)A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right),\ \text{if}\ A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\geq 0,\\ (1-\epsilon)A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right),\ \text{if}\ A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)<0,\end{cases}$
@@
Fig. 5 illustrates the impact of $A$ on the $J^{CLIP}$ function. Fig. (5a) and Fig. (5b) correspond to cases where $A\geq 0$ and $A<0$, respectively. When $A\geq 0$, the state-action pair is favorable, and we aim to increase its probability; that is, the larger $p_{\boldsymbol{\theta}}$ is, the better. However, the maximum value of $r_{t}(\boldsymbol{\theta})$ cannot exceed $1+\epsilon$, meaning the ratio of $p_{\boldsymbol{\theta}}$ to $p_{\boldsymbol{\theta^{\prime}}}$ should not exceed $1+\epsilon$. When $A<0$, the state-action pair is unfavorable, and we aim to reduce $p_{\boldsymbol{\theta}}$, but the ratio of $p_{\boldsymbol{\theta}}$ to $p_{\boldsymbol{\theta^{\prime}}}$ should not fall below $1-\epsilon$. By using the CLIP function, the difference between $p_{\boldsymbol{\theta}}$ and $p_{\boldsymbol{\theta^{\prime}}}$ is moderate, which helps improve the model’s convergence speed and performance. The loss function of the critic network aims to minimize the discrepancy between the state-value function and the cumulative reward and is expressed as
$L_{t}^{\mathrm{CL}}=\left\|{V}_{t}^{\boldsymbol{\theta^{\prime}}}\left(\boldsymbol{s}_{t}\right)-R_{t}(\tau)\right\|^{2}.$ (25)
To enhance the stochasticity of the policy, the final loss function for the MAPPO algorithm is formulated as
$L_{t}^{\mathrm{CL}+\operatorname{CLIP}+S}(\boldsymbol{\theta})=\mathbb{E}_{t}\left[L_{t}^{\mathrm{CLIP}}-b_{1}L_{t}^{\mathrm{CL}}+b_{2}H(\pi_{\boldsymbol{\theta}}\left(\cdot\mid\boldsymbol{s}_{t}\right))\right],$ (26)
@@
$\text{arg}\max_{\boldsymbol{\theta}}\hat{\mathbb{E}}_{t}\left[L_{t}^{CLIP}+b_{2}H(\pi_{\boldsymbol{\theta}}\left(\cdot\mid\boldsymbol{s}_{t}\right))\right],$ (27)
@@
Initialization: Initialize the policy newtork with parameters $\boldsymbol{\theta}$, the demonstration policy network with parameters $\boldsymbol{\theta^{\prime}}$ and the critic network with parameters $\boldsymbol{\phi}$. Initialize the observation space $o_{n}$ and the action space $a_{n}$.

### v3

@@ III-B Transmission Model
$t_{n}=\frac{\rho_{n}s_{n}\left(\mu_{n}\right)^{p}}{R_{n}},$ (2)
where $s_{n}$ is the normalized transmission data from UEn to ES, and $p>0$ is a constant parameter related to the specific task. $\rho_{n}$ is an offload indicator, where $\rho_{n}=1$ represents offloading tasks to ES, and $\rho_{n}=0$ denotes local computation. Compared to the amount of data required for task offloading, the computation results typically involve a much smaller data size [96, 97], therefore, we only consider the upload rate and latency.
@@
TABLE II: Notations and Symbols
Notations Explanations
@@
$\lambda,\beta,\eta$ The slope of the logistic function
III-C Computation Model
@@
[…]
where $q>0$ is a constant parameter relevant to the specific task, $l_{n}^{U}$ is the GPU consumption of the current task computed by UEn, and $l_{n}^{SE}$ is the basic GPU consumption of semantic extraction. $l_{n}^{U}$ and $l_{n}^{SE}$ are related to the corresponding task $\mathcal{M}$. $k>0$ is a constant parameter. $c_{n}=\alpha_{ma}N_{u}f_{n}$ is the GPU computation capability of UEn. $\alpha_{ma}=2$ corresponds to the floating-point operations (FLOPs) per cycle per core (FLOPs/cycle/core), where each instruction executes a fused multiply-add (FMA) operation. $N_{u}$ represents the number of compute unified device architecture (CUDA) cores available on UEn. $f_{n}$ denotes the GPU clock frequency of UEn. When tasks are offloaded to ES, the computation time of tasks on the ES can be expressed as
@@
III-D Energy Consumption Model
@@
$E^{S}=\kappa^{S}t^{S}\left(f_{s}\right)^{3},$ (7)
where $\kappa^{S}$ is the energy coefficient of ES. Therefore, the energy consumption of current task $q_{n}^{\mathcal{M}}$ can be expressed as
@@
III-E Multi-Task Offloading Design and Problem Formulation
@@
$\varepsilon_{n}^{\mathcal{M}}=f^{\mathcal{M}}\left(\mu_{n},\nu_{n}\right),$ (9)
where $\varepsilon_{n}^{\mathcal{M}}$ is the accuracy of task $\mathcal{M}$, and $f^{\mathcal{M}}(\cdot)$ is the pretrained semantic transmission model for task $\mathcal{M}$, and $\nu_{n}$ is SNR. For a given fixed SNR, the task performance is modeled as a power-law function of $\mu_{n}$. Once the model training is completed, we can directly calculate the number of FLOPs required by the UE’s semantic encoder to determine the basic GPU consumption $l_{n}^{SE}$ for semantic extraction. Similarly, we compute the number of FLOPs required by the ES’s semantic decoder to determine the GPU consumption $l_{n}^{S}$ for the ES. To evaluate the performance of offloading tasks for different UEs, it is essential to define a unified QoE standard.
@@
The performance of task execution is associated with computation time, energy consumption, and task accuracy. However, defining a fair and reasonable QoE is challenging. When tasks of different types are unevenly distributed among different UEs, directly calculating the weighted sum of energy consumption, computation time, and task accuracy is unfair. Typically, text tasks have lower computation time and energy consumption than image tasks, while image tasks have lower computation time and energy consumption than VQA tasks. Additionally, the accuracy of different tasks is usually incomparable. Inspired by [98, 99, 26], we propose a QoE calculation method that can fairly compare task execution performance across different UEs:
@@
where $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ are the weights of computation time, energy consumption, and task accuracy, respectively. The sum of $\omega_{t}$, $\omega_{e}$ and $\omega_{a}$ are equal to 1. $Q_{n}$ is the task length of UEn. $t^{l}_{n}=l_{n}^{U}/c_{n}$ denotes the task execution delay by UEn. As shown in Fig. 3, we use the logistic/sigmoid function and its symmetrical counterpart for QoE calculations. Corresponding latency score $G^{t}_{\mathcal{M}}$ and energy consumption score $G^{e}_{\mathcal{M}}$ are calculated using the symmetric logistic function, while accuracy score $G^{a}_{\mathcal{M}}$ is measured with the logistic function. $\lambda,\beta,\eta$ are the slope of the logistic function. We chose the logistic function for normalizing because it maps raw performance measures into a normalized interval $(0,1)$, which is suitable for probability-like interpretations and ensures stability when combining terms. Moreover, the S-shaped curve of the logistic function effectively captures the diminishing marginal utility of QoE. For example, when network resources are already abundant, further increases in resources lead to only marginal improvements in QoE. The offset $x_{0}$ represents local execution latency, energy consumption, or accuracy. Specifically, $E_{n}^{l}=\kappa^{U}t_{n}^{l}(f_{n})^{3}$ denotes the task execution energy consumption by UEn. $\varepsilon_{n}^{l}$ denotes the task execution accuracy by UEn. That is, we normalize the QoE using the latency, energy consumption, and accuracy of tasks executed locally by the UE.
@@
(P0) $\underset{\{\boldsymbol{\rho},\boldsymbol{p},\boldsymbol{f},\boldsymbol{x},\boldsymbol{\mu}\}}{\operatorname{maximize}}\quad\sum_{n=1}^{N}\mathrm{QoE}_{n}$ (11a)
@@
$\sum_{k=1}^{\mathcal{K}}x_{k,n}\leq\rho_{n},\quad\forall n\in\mathcal{N},$ (11d)
$\sum_{n=1}^{\mathcal{N}}x_{k,n}\leq 1,\quad\forall k\in\mathcal{K},$ (11e)
@@
where $\{\boldsymbol{\rho},\boldsymbol{p},\boldsymbol{f},\boldsymbol{x},\boldsymbol{\mu}\}=\{\rho_{n},p_{n},f_{n},x_{k,n},\mu_{n}\},\forall{n}\in\mathcal{N},\forall{k}\in\mathcal{K}$. $\boldsymbol{x}\in\mathbb{R}^{K\times N}$ denote the collection of all sub-channel allocation indicators $x_{k,n}$. $x_{k,n}=1$ indicates UEn is assigned to sub-channel $k$, $x_{k,n}=0$ otherwise. In (P0), (11b) and (11c) represent the offloading choice and channel selection constraints, respectively. Constraints (11d) and (11e) indicate that each UE is assigned to a maximum of one sub-channel and each sub-channel is occupied by a maximum of one UE. $p_{max}$ in (11g) is the maximum transmit power constraint, $f_{max}$ in (11h) is the maximum GPU frequency of UEn, $t_{max}$ in (11i) is the maximum execute latency constraint, $E_{n}^{max}$ in (11j) is the battery capacity of UEn, and $\varepsilon_{min}$ in (11k) is the minimum task accuracy requirement, which is dependent on the task type $\mathcal{M}$ of the current task.
IV Multi-Agent Proximal Policy Optimization for Resource Allocation
The binary offloading variable in Constraint (11b) introduces non-convexity, and the QoE is characterized by a non-convex logistic function, rendering the problem P0 a non-convex MINLP problem. To solve the non-convex MINLP problem P0, we propose an advanced semantic-aware MAPPO algorithm, i.e., an actor-critic-based reinforcement learning algorithm, which is suitable for both discrete and continuous action spaces. PPO-based DRL algorithms are known for their stable and reliable updates due to the clipped surrogate objective, which constrains large policy changes during training. This stability is particularly important in communication systems. By contrast, the asynchronous advantage actor-critic (A3C) algorithm uses asynchronous parallel updates without explicit trust-region constraints, which can lead to high variance and unstable learning in multi-agent settings. The multi-agent deep deterministic policy gradient (MADDPG) algorithm, while designed for multi-agent control, relies on centralized action–value critics and deterministic policies that can be sensitive to hyperparameters and suffer from exploration issues in high-dimensional observation spaces.
To utilize the semantic-aware MAPPO algorithm, we reformulate problem P0 as an MDP and detail the training process of the MAPPO.
IV-A Problem Reformulation Based on MDP
@@
A 4-tuple MDP $\langle\mathcal{S},\mathcal{A},\mathcal{R},\gamma\rangle$ is used to describe the interaction between the ES and the wireless network, where $\mathcal{S}$, $\mathcal{A}$, $\mathcal{R}$ and $\gamma\in[0,1)$ denote the state space, action space, reward function, and the discount factor, respectively. As agents, UEs need to collect observable states to input into the policy network and obtain their actions. After executing actions, UEs update their observable states and the policy network based on corresponding rewards to improve the output of actions. Specifically, the observable state of UEn can be represented as $o_{n}=\left\{g_{k,n},l_{n}^{U},l_{n}^{S},t_{n}^{U},E_{n}^{U},\varepsilon_{n}^{\mathcal{M}},t_{max},E_{n}^{max},\varepsilon_{min}\right\}\in\mathcal{S}$. The action space of UEn refers to the variables that need to be optimized, which can be expressed as $a_{n}=\left\{\rho_{n},p_{n},f_{n},\mu_{n},x_{k,n}\right\}\in\mathcal{A}$. UEs need to continuously optimize offloading strategies, transmit power, local computation frequency, semantic compression ratio, and channel assignment to purchase better QoE. Drawing from the optimization problem P0, it is evident that the reward function needs to focus on the summation QoE of all UEs. In addition to QoE, the reward function includes normalized penalty terms for latency, energy consumption, and task accuracy, expressed as
$r_{n}=\text{QoE}_{n}-\left[\frac{t^{\mathcal{M}}_{n}-t_{max}}{t_{max}}\right]_{+}$- $\left[\frac{E_{n}^{\mathcal{M}}-E_{n}^{max}}{E_{n}^{max}}\right]_{+}$ (12)
- $\left[\frac{\varepsilon_{min}-\varepsilon_{n}^{\mathcal{M}}}{\varepsilon_{min}}\right]_{+},$
@@
where $\left[\cdot\right]_{+}=\max(\cdot,0)$. Therefore, the total reward at the $t$th training timestep is defined as $r_{t}=\sum^{N}_{n=1}r_{n}$. The cumulative long-term reward accounts for the impact of the current state and action on the expected future rewards:
@@
IV-B Semantic-aware MAPPO Algorithm
@@
$J^{\boldsymbol{\theta}^{\prime}}(\boldsymbol{\theta})=\mathbb{E}_{\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\sim\pi_{\boldsymbol{\theta}^{\prime}}}\left[\frac{p_{\boldsymbol{\theta}}\left(\boldsymbol{a}_{t}\mid\boldsymbol{s}_{t}\right)}{p_{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{a}_{t}\mid\boldsymbol{s}_{t}\right)}A^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\right],$ (17)
@@
$A^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=Q^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)-V^{\boldsymbol{\phi}}\left(\boldsymbol{s}_{t}\right),$ (18)
where $\boldsymbol{\phi}$ is the parameters of critic network. The state-action value function $Q^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)$ is represented as
$Q^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=\mathbb{E}_{\boldsymbol{a}_{t}\sim\pi_{\boldsymbol{\theta^{\prime}}},\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}\right],$ (19)
@@
$V^{\boldsymbol{\phi}}\left(\boldsymbol{s}_{t}\right)=\mathbb{E}_{\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}\right].$ (20)
It is noted that $Q^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)$ can also be expressed by temporal difference (TD) form as $Q^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=r_{t}+\gamma V^{\boldsymbol{\phi}}\left(\boldsymbol{s}_{t+1}\right)$. Therefore, (18) can be reformulated as
$A^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=r_{t}+\gamma V^{\boldsymbol{\phi}}\left(\boldsymbol{s}_{t+1}\right)-V^{\boldsymbol{\phi}}\left(\boldsymbol{s}_{t}\right).$ (21)
@@
$A^{\mathrm{GAE}}(\boldsymbol{s}_{t},\boldsymbol{a}_{t})=\sum_{l=0}^{T-t}(\lambda\gamma)^{l}\delta_{t+l},$ (22)
where $\delta_{t}=A^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)$. $\lambda\in(0,1]$ is a discount factor. Note that (22) aims to balance the variance and bias of the advantage estimates by leveraging information from multiple future steps while discounting their influence.
@@
If there are no constraints on the policies $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}^{\prime}}$, significant fluctuations may occur during policy updates. This is due to the limited sampling, which fails to accurately represent the probability distributions described in Lemma 1. To constrain the divergence between policies $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}^{\prime}}$, $J^{\mathrm{CLIP}}(\boldsymbol{\theta})$ is employed. It is defined as
$J^{\mathrm{CLIP}}(\boldsymbol{\theta})=\mathbb{E}_{t}\Big[\min\Big(r_{t}(\boldsymbol{\theta})A^{\mathrm{GAE}}(\boldsymbol{s}_{t},\boldsymbol{a}_{t}),$ (23)
$\qquad g\!\left(\epsilon,A^{\mathrm{GAE}}(\boldsymbol{s}_{t},\boldsymbol{a}_{t})\right)\Big)\Big],$
@@
$g(\epsilon,$ $A^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right))=$ (24)
$(1ϵ)AGAE(𝒔t,𝒂t),ifAGAE(𝒔t,𝒂t)<0,\displaystyle$
@@
Fig. 5 illustrates the impact of $A$ on the $J^{\mathrm{CLIP}}$ function. Fig. 5a and Fig. 5b correspond to cases where $A\geq 0$ and $A<0$, respectively. When $A\geq 0$, the state-action pair is favorable, and we aim to increase its probability; that is, the larger $p_{\boldsymbol{\theta}}$ is, the better. However, the maximum value of $r_{t}(\boldsymbol{\theta})$ cannot exceed $1+\epsilon$, meaning the ratio of $p_{\boldsymbol{\theta}}$ to $p_{\boldsymbol{\theta^{\prime}}}$ should not exceed $1+\epsilon$. When $A<0$, the state-action pair is unfavorable, and we aim to reduce $p_{\boldsymbol{\theta}}$, but the ratio of $p_{\boldsymbol{\theta}}$ to $p_{\boldsymbol{\theta^{\prime}}}$ should not fall below $1-\epsilon$. By using the CLIP function, the difference between $p_{\boldsymbol{\theta}}$ and $p_{\boldsymbol{\theta^{\prime}}}$ is moderate, which helps improve the model’s convergence speed and performance. The loss function of the critic network aims to minimize the discrepancy between the state-value function and the cumulative reward and is expressed as
$L_{t}^{\mathrm{CL}}=\left\|{V}^{\boldsymbol{\phi}}\left(\boldsymbol{s}_{t}\right)-R_{t}(\tau)\right\|^{2}.$ (25)
To enhance the stochasticity of the policy, the ultimate optimization object for the MAPPO algorithm is formulated as
$L_{t}^{\mathrm{CL+CLIP+S}}(\theta)=\mathbb{E}_{t}\Big[$ $J_{t}^{\mathrm{CLIP}}-b_{1}L_{t}^{\mathrm{CL}}$ (26)
$+b_{2}H\!\left(\pi_{\theta}(\cdot\mid s_{t})\right)\Big],$
@@
$\text{arg}\max_{\boldsymbol{\theta}}\hat{\mathbb{E}}_{t}\left[J_{t}^{\mathrm{CLIP}}+b_{2}H(\pi_{\boldsymbol{\theta}}\left(\cdot\mid\boldsymbol{s}_{t}\right))\right],$ (27)
@@
Initialization: Initialize the policy network with parameters $\boldsymbol{\theta}$, the sampling policy network with parameters $\boldsymbol{\theta^{\prime}}$ and the critic network with parameters $\boldsymbol{\phi}$. Initialize the observation space $o_{n}$ and the action space $a_{n}$.